# Homological algebra of pro-Lie Polish abelian groups

### v1

@@ 2.6. The left heart of thick categories of Polish abelian groups
Let $\mathbf{PAb}$ be the category of Polish abelian groups and continuous group homomorphism. This is a quasi-abelian category [25, Lemma 6.1] An explicit description of LH$\left(\mathbf{PAb}\right)$ was given in [25] in terms of groups with a Polish cover. An abelian group with a Polish cover is a group $G$ explicitly presented as a quotient $\hat{G}/N$, where $\hat{G}$ is a Polish abelian group and $N\subseteq\hat{G}$ is a Polishable subgroup of $\hat{G}$. This means that $N$ is a Polish group with respect to some Polish topology such that the inclusion $N\rightarrow\hat{G}$ is continuous. The Polish topology on $N$ is in general not the subspace topology inherited from $\hat{G}$ (unless $N$ is closed in $\hat{G}$). Nonetheless, the topology on $N$ is in some sense induced by the topology on $\hat{G}$, as it is the unique Polish topology on $N$ whose open sets are Borel in $\hat{G}$. Every Polish abelian group $G$ can be identified with the group with a Polish cover $\hat{G}/N$ where $G=\hat{G}$ and $N$ is the trivial subgroup of $g$. A group homomorphism $\varphi:G\rightarrow H$ between groups with a Polish cover $G=\hat{G}/N$ and $H=\hat{H}/M$ is Borel-definable if has a lift to a Borel function $f:\hat{G}\rightarrow\hat{H}$, such that $\varphi\left(x+N\right)=f\left(x\right)+M$ for every $x\in\hat{G}$.
The category $\mathrm{LH}\left(\mathbf{PAb}\right)$ is (equivalent to) the category of groups with a Polish cover and Borel-definable group homomorphisms [25, Theorem 6.2]. More generally, if $\mathcal{B}$ is a thick subcategory of $\mathbf{PAb}$, then $\mathrm{LH}\left(\mathcal{B}\right)$ is (equivalent to) the full subcategory of the category of abelian groups with a Polish cover spanned by the groups with a Polish cover of the form $\hat{G}/N$ where both $\hat{G}$ and $N$ belong to $\mathcal{B}$ [25, Theorem 6.13 and Proposition 6.15].
@@ 3.1. Borel cocycles for Polish groups
Let $G,H$ be abelian Polish groups. A (symmetric,) $2$-cocycle on $G$ with values in $H$ is a function $c:G\times G\rightarrow H$ satisfying the following identities for all $x,y,z\in G$:
@@
We let $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)$ be the group whose elements are the isomorphism classes of abelian Polish group extensions $A\rightarrow X\rightarrow C$, where the group operation is induced by Baer sum of extensions and the trivial element is the class of split extensions. A short exact sequence $A\rightarrow X\rightarrow C$ in $\mathbf{PAb}$ yields an element of $\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ as follows. Pick a Borel right inverse $t:C\rightarrow X$ for the map $X\rightarrow C$, which exists by [23, Theorem 12.17]. Identifying $A$ with a closed subgroup of $X$, we have that for every $x,y\in C$, $\kappa\left(x,y\right):=t\left(x+y\right)-t\left(x\right)-t\left(y\right)$ belongs to $A$. This defines a Borel $2$-cocycle on $C$ with values in $A$, whose corresponding element of $\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ is independent on the choice of $t$. This assignment defines an injective group homomorphism $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$.
@@
Suppose that $C$ and $A$ are locally compact Polish group. Then the group homomorphism $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ is an isomorphism.
@@
We also endow $A\times C$ with the product Borel structure and the Borel measure defined as the product of Haar measures on $A$ and $G$. Then we have that this is a $\sigma$-finite invariant Borel measure, and so by [27, Theorem 1] there exists a unique locally compact Polish group topology on $X$ that is compatible with its Borel structure and has the given measure as a Haar measure. When endowed with this topology, the canonical maps $A\rightarrow X\rightarrow C$ give a short exact sequence in $\mathbf{PAb}$. The image of the corresponding element in $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)$ under the homomorphism $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ is equal to the element of $\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ represented by $\kappa$. ∎
@@
Given an short exact sequence $A\rightarrow B\rightarrow C$ with corresponding cocycle $\kappa$ as above, and abelian Polish groups $X$ and $Y$, we define the corresponding boundary homomorphisms $\mathrm{Hom}\left(X,C\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)$, $\varphi\mapsto\kappa\circ\left(\varphi\times\varphi\right)+\mathrm{B}^{1}\left(X,A\right)$ and Hom$\left(A,Y\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(C,Y\right)$, $\psi\mapsto\psi\circ\kappa+\mathrm{B}^{1}\left(C,Y\right)$. The same proof as in the case of discrete groups gives the following lemma; see [15, Chapter IX].
@@
    If $X$ is non-Archimedean, then we have an exact sequence of abelian groups:
@@
Let us fix a Borel right inverse $\sigma:C\rightarrow B$ for $\pi$, and let $\kappa=\delta\sigma:C\times C\rightarrow A$. We identify $A$ with a closed subgroup of $B$.
(1) and (2): It is clear that the image of $\mathrm{Hom}\left(X,C\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)$ is contained in the kernel of $\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)$. We prove the converse implication. Suppose that $c:X\times X\rightarrow A$ is a Borel cocycle such that there exists a function $t:X\rightarrow B$ such that $c=\delta t$. Then we have that $\pi t:X\rightarrow C$ is a group homomorphism. Then $\kappa\circ\left(\pi t\times\pi t\right)$ is a cocycle cohomologous to $c$. Define $g:X\rightarrow A$, $x\mapsto\sigma\pi t\left(x\right)-t\left(x\right)$. For $x,y\in X$,
@@
This concludes the proof that the image of $\mathrm{Hom}\left(X,C\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)$ is equal to the kernel of $\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)$
We now prove that the kernel of $\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,C\right)$ is contained in the image of $\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)$. Suppose that $c:X\times X\rightarrow B$ is a Borel cocycle such that there is a Borel function $f:X\rightarrow C$ such that $\pi c=\delta f$. Then we have that $c_{0}:=c-\delta\left(\sigma f\right)$ is a Borel cocycle $X\times X\rightarrow A$ such that $c=c_{0}+\delta\left(\sigma f\right)$ and hence $c,c_{0}$ are cohomologous as Borel cocycles $X\times X\rightarrow B$. This shows that the kernel of $\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,C\right)$ is contained in the image of $\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)$. If $c$ is continuous, then $f$ is continuous, and hence $c_{0}$ is continuous as well. This shows that the kernel of $\mathrm{Ext}_{\mathrm{Yon}}\left(X,B\right)\rightarrow\mathrm{Ext}_{\mathrm{Yon}}\left(X,C\right)$ is contained in the image of $\mathrm{Ext}_{\mathrm{Yon}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{Yon}}\left(X,B\right)$ when $X$ is non-Archimedean.
@@
$f\left(b\right)+f\left(-b\right)=-c\left(b,-b\right)\text{.}$
Then we have that $\varphi:=f|_{A}:A\rightarrow Y$ is a continuous group homomorphism. We have that $\varphi\circ\kappa$ is a Borel cocycle cohomologous to $c$. Define $g:=f\sigma:C\rightarrow Y$. For $x,y\in C$ we have that
@@
is a push-out. This yields a short exact sequence $Y\rightarrow H/\xi\left(A\right)\rightarrow C$ that represents an element of $\mathrm{Ext}\left(C,Y\right)$ whose image in $\mathrm{Ext}_{\mathrm{Yon}}\left(B,Y\right)$ is the element represented by $Y\rightarrow H\rightarrow B$. ∎
@@
$\mathrm{Ext}\left(G,H\right)\cong\prod_{n\in\alpha}\mathrm{Ext}\left(G,H_{n}\right)\text{.}$
Furthermore, if $\alpha<\omega$ then, since $\mathrm{Ext}\left(-,G\right)$ is an additive functor, we also have
$\mathrm{Ext}(H,G)\cong\prod_{n\in\alpha}\mathrm{Ext}\left(H_{n},G\right)$
@@ 3.2. Preliminaries on pro-Lie groups
In this section, we recall fundamental facts about the theory of pro-Lie groups as can be found in [19]. We will only consider abelian Polish groups. The category $\mathbf{PAb}$ of abelian Polish groups and continuous group homomorphisms is a quasi-abelian category [25, Theorem 6.2]. Locally compact Polish abelian groups form a thick subcategory $\mathbf{LCPAb}$ of $\mathbf{PAb}$ [25, Theorem 6.17].
An abelian Polish group is a Lie group if and only if it is of the form $V\oplus T\oplus D$ where $D$ is countable discrete, $V$ is a finite-dimensional vector group (isomorphic to $\mathbb{R}^{n}$ for some $n\in\omega$), and $T$ is a finite-dimensional torus (isomorphic to $\mathbb{T}^{d}$ for some $d\in\omega$); see [20, Exercise E5.18 following]. Lie groups form a thick subcategory $\mathbf{LiePAb}$ of the category of locally compact abelian Polish groups; see [25, Theorem 6.17].
A Polish abelian group $G$ has no small subgroups if and only if it has a zero neighborhood $U$ such that for every subgroup $N$ of $G$ contained in $U$, one has that $N=\left\{0\right\}$ [29]. When $G$ is locally compact, this is equivalent to the assertion that the Pontryagin dual $G^{\vee}$ is compactly generated, as well as to the assertion that $G$ is a Lie group [29, Theorem 2.4 and Corollary 1].
An abelian Polish group is non-Archimedean if it has a basis of zero neighborhoods consisting of subgroups. This is equivalent to the assertion that $A$ is isomorphic to an inverse limit of countable groups. Non-Archimedean abelian Polish groups form a fully exact quasi-abelian subcategory of the category of pro-Lie Polish abelian groups; see [25, Theorem 6.17].
@@
An abelian Polish group $G$ is a pro-Lie group if every zero neighborhood of $G$ contains a closed subgroup $N$ such that $G/N$ is a Lie group. This implies that $\mathcal{N}\left(G\right)$ is closed under intersections, and hence a filter basis [19, page 148]. It also implies that $G\cong\mathrm{\mathrm{lim}}_{N\in\mathcal{N}\left(G\right)}G/N$. For a decreasing chain $\left(N_{k}\right)_{k\in\omega}$ in $\mathcal{N}\left(G\right)$, we say that $N_{k}\rightarrow 0$ if for every zero neighborhood $U$ in $G$, $N_{k}$ is contained in $U$ eventually.
@@
Suppose that $N_{k}\rightarrow 0$. For $M\in\mathcal{N}\left(G\right)$, we have that $G/M$ is an a Lie group. Thus, it has a zero neighborhood $U$ that does not contain any subgroup. Since $N_{k}\rightarrow 0$, we have that there exists $k\in\omega$ such that $\pi\left(N_{k}\right)\subseteq U$, where $\pi:G\rightarrow G/M$ is the quotient map. This implies that $\pi\left(N_{k}\right)=\left\{0\right\}$ and $N_{k}\subseteq M$. The converse implication is obvious. The second assertion follows from the fact that $G\cong\mathrm{\mathrm{lim}}_{N\in\mathcal{N}\left(G\right)}G/N$. ∎
It is proved in [19, Theorem 3.39] that a Polish abelian group is pro-Lie if and only if it is the inverse limit of an inverse system of Polish abelian Lie groups.
A Polish group $G$ is almost connected if $G/G_{0}$ is compact, where $G_{0}$ is the connected component of the trivial element of $G$ (which is a closed subgroup of $G$) [19, Definition 5.6]. If $G$ is an abelian pro-Lie group, then $G/G_{0}$ is non-Archimedean [19, Theorem 5.20(iii)].
The class of abelian Polish pro-Lie groups contains all locally compact abelian Polish groups [19, Example 5.1] and all almost connected abelian Polish groups [37, 38], and is closed within the category of Polish groups under the following operations [19, Chapter 3]:
@@
We also say that an abelian Polish pro-Lie group is vector-free if it has no nonzero closed subgroups that are vector groups. By [19, Theorem 5.19] we have the following:
@@
An abelian Polish pro-Lie group $G$ admits a closed subgroup $V$, called maximal vector subgroup (or vector group complement), that is a vector group, and such that $H:=G/V$ is vector-free; see [19, Theorem 5.20]. In this case, by Lemma 3.7 we have that $G\cong V\oplus H$. Weil’s Lemma for pro-Lie groups asserts the following; see [19, Theorem 5.3].
@@
Let $G$ be an abelian Polish pro-Lie group and let $E$ be either $\mathbb{Z}$ or $\mathbb{R}$. If $f:E\rightarrow G$ is a continuous homomorphism, then exactly one of the following alternatives holds: either the image of $f$ has compact closure, or $f$ is injective.
Let $G$ be an abelian Polish group. Define $\mathrm{comp}\left(G\right)$ to be the subgroup of $g\in G$ such that the subgroup generated by $g$ has compact closure. If $G$ is an abelian Polish pro-Lie group, then $\mathrm{comp}\left(G\right)$ is a closed subgroup of $G$ [19, Theorem 5.5]. One says that $G$ is elementwise compact if $G=\mathrm{comp}\left(G\right)$ and compact-free if $\mathrm{comp}\left(G\right)=\left\{0\right\}$; see [19, Definition 5.4]. We have that $G/\mathrm{comp}\left(G\right)\cong V\oplus S$ where $V$ is a maximal vector subgroup and $S$ is non-Archimedean and compact-free [19, Theorem 5.20(iv)]. By [19, Proposition 5.43], we have the following:
@@
The following result is established in [19, Theorem 4.1].
@@
We regard pro-Lie Polish abelian groups as a full subcategory $\mathbf{proLiePAb}$ of the quasi-abelian category $\mathbf{PAb}$ of Polish abelian groups. We will soon prove that $\mathbf{proLiePAb}$ is in fact a thick subcategory of $\mathbf{PAb}$. It follows from Proposition 3.12 that $\mathbf{proLiePAb}$ is isomorphic to the category $\mathrm{pro}_{\omega}\left(\mathbf{LieAb}\right)$ of towers of Lie Polish abelian groups. This is the full subcategory of the pro-category $\mathrm{pro}\left(\mathbf{LieAb}\right)$ spanned by the inverse systems indexed by the ordered set $\omega$.
When $G$ is locally compact and $H$ is pro-Lie Polish, we have that $\mathrm{Hom}\left(G,H\right)$ is a Polish abelian group endowed with the compact-open topology, and the isomorphism
@@
By Lemma 3.13, the vector groups are precisely the pro-Lie Polish abelian groups that are Polish $\mathbb{R}$-vector spaces. These are also called weakly complete topological Polish $\mathbb{R}$-vector spaces in [19, Proposition 5.43]. We have that if $V,W$ are vector groups, and $\varphi:V\rightarrow W$ is a continuous homomorphism, then it is $\mathbb{R}$-linear and its image is a closed subspace of $W$ which is a direct summand [19, Theorem A2.12].

### v2

@@ 2.6. The left heart of thick categories of Polish abelian groups
Let $\mathbf{PAb}$ be the category of Polish abelian groups and continuous group homomorphisms. This is a quasi-abelian category [26, Lemma 6.1]. An explicit description of LH$\left(\mathbf{PAb}\right)$ was given in [26] in terms of groups with a Polish cover. An abelian group with a Polish cover is a group $G$ explicitly presented as a quotient $\hat{G}/N$, where $\hat{G}$ is a Polish abelian group and $N\subseteq\hat{G}$ is a Polishable subgroup of $\hat{G}$. This means that $N$ is a Polish group with respect to some Polish topology such that the inclusion $N\rightarrow\hat{G}$ is continuous. The Polish topology on $N$ is in general not the subspace topology inherited from $\hat{G}$ (unless $N$ is closed in $\hat{G}$). Nonetheless, the topology on $N$ is in some sense induced by the topology on $\hat{G}$, as it is the unique Polish topology on $N$ whose open sets are Borel in $\hat{G}$. Every Polish abelian group $G$ can be identified with the group with a Polish cover $\hat{G}/N$ where $G=\hat{G}$ and $N$ is the trivial subgroup of $G$. A group homomorphism $\varphi:G\rightarrow H$ between groups with a Polish cover $G=\hat{G}/N$ and $H=\hat{H}/M$ is Borel-definable if has a lift to a Borel function $f:\hat{G}\rightarrow\hat{H}$, such that $\varphi\left(x+N\right)=f\left(x\right)+M$ for every $x\in\hat{G}$.
The category $\mathrm{LH}\left(\mathbf{PAb}\right)$ is (equivalent to) the category of groups with a Polish cover and Borel-definable group homomorphisms [26, Theorem 6.2]. More generally, if $\mathcal{B}$ is a thick subcategory of $\mathbf{PAb}$, then $\mathrm{LH}\left(\mathcal{B}\right)$ is (equivalent to) the full subcategory of the category of abelian groups with a Polish cover spanned by groups with a $\mathcal{B}$-cover, which are the groups with a Polish cover of the form $\hat{G}/N$ where both $\hat{G}$ and $N$ belong to $\mathcal{B}$ [26, Theorem 6.13 and Proposition 6.15]. As it is remarked therein, a Borel-definable homomorphism between groups with a $\mathcal{B}$-cover is an isomorphism in $\mathrm{LH}\left(\mathcal{B}\right)$ if and only if it is a bijection.
An abelian Polish group is non-Archimedean if it has a basis of zero neighborhoods consisting of subgroups. This is equivalent to the assertion that $A$ is isomorphic to an inverse limit of countable groups. Non-Archimedean Polish abelian groups form a thick subcategory of the category of Polish abelian groups; see [26, Theorem 6.17].
@@ 3.1. Borel cocycles for Polish groups
Let $G,H$ be abelian Polish groups. A (symmetric) $2$-cocycle on $G$ with values in $H$ is a function $c:G\times G\rightarrow H$ satisfying the following identities for all $x,y,z\in G$:
@@
We let $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)$ be the group whose elements are the isomorphism classes of abelian Polish group extensions $A\rightarrow X\rightarrow C$, where the group operation is induced by Baer sum of extensions and the trivial element is the class of split extensions. A short exact sequence $A\rightarrow X\rightarrow C$ in $\mathbf{PAb}$ yields an element of $\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ as follows. Pick a Borel right inverse $t:C\rightarrow X$ for the map $X\rightarrow C$, which exists by [24, Theorem 12.17]. Identifying $A$ with a closed subgroup of $X$, we have that for every $x,y\in C$, $\kappa\left(x,y\right):=t\left(x+y\right)-t\left(x\right)-t\left(y\right)$ belongs to $A$. This defines a Borel $2$-cocycle on $C$ with values in $A$, whose corresponding element of $\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ is independent on the choice of $t$. This assignment defines an injective group homomorphism $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$.
@@
Suppose that $C$ and $A$ are locally compact Polish abelian groups. Then the group homomorphism $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ is an isomorphism.
@@
We also endow $A\times C$ with the product Borel structure and the Borel measure defined as the product of Haar measures on $A$ and $G$. Then we have that this is a $\sigma$-finite invariant Borel measure, and so by [28, Theorem 1] there exists a unique locally compact Polish group topology on $X$ that is compatible with its Borel structure and has the given measure as a Haar measure. When endowed with this topology, the canonical maps $A\rightarrow X\rightarrow C$ give a short exact sequence in $\mathbf{PAb}$. The image of the corresponding element in $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)$ under the homomorphism $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ is equal to the element of $\mathrm{Ext}_{\mathrm{c}}\left(C,A\right)$ represented by $\kappa$. ∎
@@
Given a short exact sequence $A\rightarrow B\rightarrow C$ with the corresponding cocycle $\kappa$ as above, and abelian Polish groups $X$ and $Y$, we define the corresponding boundary homomorphisms $\mathrm{Hom}\left(X,C\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)$, $\varphi\mapsto\kappa\circ\left(\varphi\times\varphi\right)+\mathrm{B}^{1}\left(X,A\right)$ and Hom$\left(A,Y\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(C,Y\right)$, $\psi\mapsto\psi\circ\kappa+\mathrm{B}^{1}\left(C,Y\right)$. The same proof as in the case of discrete groups gives the following lemma; see [15, Chapter IX]. When $X,C$ are non-Archimedean, one can choose $\kappa$ to be continuous, whence the boundary homomorphism takes values in $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)$.
@@
    If $X,C$ are non-Archimedean, then we have an exact sequence of abelian groups:
@@
Let us fix a Borel right inverse $\sigma:C\rightarrow B$ for $\pi$, and let $\kappa=\delta\sigma:C\times C\rightarrow A$. We identify $A$ with a closed subgroup of $B$. By [4, Proposition 4.6], if $C$ is non-Archimedean, then one can choose $\sigma$ to be continuous.
(1) and (2): It is clear that the image of $\mathrm{Hom}\left(X,C\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)$ is contained in the kernel of $\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)$. We prove the converse implication. Suppose that $c:X\times X\rightarrow A$ is a Borel cocycle such that there exists a function $t:X\rightarrow B$ such that $c=\delta t$. Then we have that $\pi t:X\rightarrow C$ is a group homomorphism. Then $\kappa\circ\left(\pi t\times\pi t\right)$ is a cocycle cohomologous to $c$, as we are about to show. Define $g:X\rightarrow A$, $x\mapsto\sigma\pi t\left(x\right)-t\left(x\right)$. For $x,y\in X$,
@@
This concludes the proof that the image of $\mathrm{Hom}\left(X,C\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)$ is equal to the kernel of $\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)$.
We now prove that the kernel of $\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,C\right)$ is contained in the image of $\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)$. Suppose that $c:X\times X\rightarrow B$ is a Borel cocycle such that there is a Borel function $f:X\rightarrow C$ such that $\pi c=\delta f$. Then we have that $c_{0}:=c-\delta\left(\sigma f\right)$ is a Borel cocycle $X\times X\rightarrow A$ such that $c=c_{0}+\delta\left(\sigma f\right)$ and hence $c,c_{0}$ are cohomologous as Borel cocycles $X\times X\rightarrow B$. This shows that the kernel of $\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,C\right)$ is contained in the image of $\mathrm{Ext}_{\mathrm{c}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{c}}\left(X,B\right)$.
When $X,C$ are non-Archimedean, $\sigma$ is continuous. If $c$ is continuous, then $f$ is continuous, and hence $c_{0}$ is continuous as well. This shows that the kernel of $\mathrm{Ext}_{\mathrm{Yon}}\left(X,B\right)\rightarrow\mathrm{Ext}_{\mathrm{Yon}}\left(X,C\right)$ is contained in the image of $\mathrm{Ext}_{\mathrm{Yon}}\left(X,A\right)\rightarrow\mathrm{Ext}_{\mathrm{Yon}}\left(X,B\right)$ when $X,C$ are non-Archimedean.
@@
$f\left(b\right)+f\left(-b\right)=-c\left(\pi\left(b\right),-\pi\left(b\right)\right)\text{.}$
Then we have that $\varphi:=f|_{A}:A\rightarrow Y$ is a continuous group homomorphism. We have that $\varphi\circ\kappa$ is a Borel cocycle cohomologous to $c$. Indeed, define $g:=f\sigma:C\rightarrow Y$. For $x,y\in C$ we have that
@@
is a push-out. This yields a short exact sequence $Y\rightarrow H/\xi\left(A\right)\rightarrow C$ that represents an element of $\mathrm{Ext}_{\mathrm{Yon}}\left(C,Y\right)$ whose image in $\mathrm{Ext}_{\mathrm{Yon}}\left(B,Y\right)$ is the element represented by $Y\rightarrow H\rightarrow B$. ∎
@@
$\mathrm{Ext}_{\mathrm{c}}\left(G,H\right)\cong\prod_{n\in\alpha}\mathrm{Ext}_{\mathrm{c}}\left(G,H_{n}\right)\text{.}$
Furthermore, if $\alpha<\omega$ then, since $\mathrm{Ext}_{\mathrm{c}}\left(-,G\right)$ is an additive functor, we also have
$\mathrm{Ext}_{\mathrm{c}}(H,G)\cong\prod_{n\in\alpha}\mathrm{Ext}_{\mathrm{c}}\left(H_{n},G\right)$
@@ 3.2. Preliminaries on pro-Lie groups
In this section, we recall fundamental facts about the theory of pro-Lie groups as can be found in [20]. We will only consider abelian Polish groups. The category $\mathbf{PAb}$ of abelian Polish groups and continuous group homomorphisms is a quasi-abelian category [26, Theorem 6.2]. Locally compact Polish abelian groups form a thick subcategory $\mathbf{LCPAb}$ of $\mathbf{PAb}$ [26, Theorem 6.17].
An abelian Polish group is a Lie group if and only if it is of the form $V\oplus T\oplus D$ where $D$ is countable discrete, $V$ is a finite-dimensional vector group (isomorphic to $\mathbb{R}^{n}$ for some $n\in\omega$), and $T$ is a finite-dimensional torus (isomorphic to $\mathbb{T}^{d}$ for some $d\in\omega$); see [21, Exercise E5.18]. Lie groups form a thick subcategory $\mathbf{LiePAb}$ of the category of locally compact abelian Polish groups; see [26, Theorem 6.17].
A Polish abelian group $G$ has no small subgroups if and only if it has a zero neighborhood $U$ such that for every subgroup $N$ of $G$ contained in $U$, one has that $N=\left\{0\right\}$ [30]. When $G$ is locally compact, this is equivalent to the assertion that the Pontryagin dual $G^{\vee}$ is compactly generated, as well as to the assertion that $G$ is a Lie group [30, Theorem 2.4 and Corollary 1].
@@
An abelian Polish group $G$ is a pro-Lie group if every zero neighborhood of $G$ contains a closed subgroup $N$ such that $G/N$ is a Lie group. This implies that $\mathcal{N}\left(G\right)$ is closed under intersections, and hence a filter basis [20, page 148]. It also implies that $G\cong\mathrm{\mathrm{lim}}_{N\in\mathcal{N}\left(G\right)}G/N$. For a decreasing chain $\left(N_{k}\right)_{k\in\omega}$ in $\mathcal{N}\left(G\right)$, we say that $N_{k}\rightarrow 0$ if for every zero neighborhood $U$ in $G$, $N_{k}$ is contained in $U$ eventually. Notice that such a sequence exists for a pro-Lie Polish abelian group (as a Polish abelian group has a countable basis of zero neighborhoods).
@@
Suppose that $N_{k}\rightarrow 0$. For $M\in\mathcal{N}\left(G\right)$, we have that $G/M$ is a Lie group. Thus, it has a zero neighborhood $U$ that does not contain any subgroup. Since $N_{k}\rightarrow 0$, we have that there exists $k\in\omega$ such that $\pi\left(N_{k}\right)\subseteq U$, where $\pi:G\rightarrow G/M$ is the quotient map. This implies that $\pi\left(N_{k}\right)=\left\{0\right\}$ and $N_{k}\subseteq M$. The converse implication is obvious. The second assertion follows from the fact that $G\cong\mathrm{\mathrm{lim}}_{N\in\mathcal{N}\left(G\right)}G/N$. ∎
It is proved in [20, Theorem 3.39] that a Polish abelian group is pro-Lie if and only if it is the inverse limit of an inverse system of Polish abelian Lie groups.
A Polish group $G$ is almost connected if $G/c\left(G\right)$ is compact, where $c\left(G\right)$ is the connected component of the trivial element of $G$ (which is a closed subgroup of $G$) [20, Definition 5.6]. If $G$ is an abelian pro-Lie group, then $G/c\left(G\right)$ is non-Archimedean [20, Theorem 5.20(iii)].
The class of abelian Polish pro-Lie groups contains all locally compact abelian Polish groups [20, Example 5.1], all non-Archimedean abelian Polish groups, and all almost connected abelian Polish groups [38, 39], and is closed within the category of Polish groups under the following operations [20, Chapter 3]:
@@
We also say that an abelian Polish pro-Lie group is vector-free if it has no nonzero closed subgroups that are vector groups. By [20, Theorem 5.19] we have the following:
@@
An abelian Polish pro-Lie group $G$ admits a closed subgroup $V$, called maximal vector subgroup (or vector group complement), that is a vector group, and such that $H:=G/V$ is vector-free; see [20, Theorem 5.20]. In this case, by Lemma 3.7 we have that $G\cong V\oplus H$. Weil’s Lemma for pro-Lie groups asserts the following; see [20, Theorem 5.3].
@@
Let $G$ be an abelian Polish pro-Lie group and let $E$ be either $\mathbb{Z}$ or $\mathbb{R}$. If $f:E\rightarrow G$ is a continuous homomorphism, then exactly one of the following alternatives holds: either the image of $f$ has compact closure, or $f$ is injective with closed image.
Let $G$ be an abelian Polish group. Define $\mathrm{comp}\left(G\right)$ to be the subgroup of $g\in G$ such that the subgroup generated by $g$ has compact closure. If $G$ is an abelian Polish pro-Lie group, then $\mathrm{comp}\left(G\right)$ is a closed subgroup of $G$ [20, Theorem 5.5]. One says that $G$ is elementwise compact if $G=\mathrm{comp}\left(G\right)$ and compact-free if $\mathrm{comp}\left(G\right)=\left\{0\right\}$; see [20, Definition 5.4]. We have that $G/\mathrm{comp}\left(G\right)\cong V\oplus S$ where $V$ is a maximal vector subgroup and $S$ is non-Archimedean and compact-free [20, Theorem 5.20(iv)]. By [20, Proposition 5.43], we have the following:
@@
The following result is established in [20, Theorem 4.1].
@@
We regard pro-Lie Polish abelian groups as a full subcategory $\mathbf{proLiePAb}$ of the quasi-abelian category $\mathbf{PAb}$ of Polish abelian groups. We will soon prove that $\mathbf{proLiePAb}$ is in fact a thick subcategory of $\mathbf{PAb}$.
When $G$ is a locally compact Polish abelian group and $H$ is a pro-Lie Polish abelian group, we have that $\mathrm{Hom}\left(G,H\right)$ is a Polish abelian group endowed with the compact-open topology [19, Exercise 1.1.6], and the isomorphism
@@
By Lemma 3.13, the vector groups are precisely the pro-Lie Polish abelian groups that are Polish $\mathbb{R}$-vector spaces. These are also called weakly complete topological Polish $\mathbb{R}$-vector spaces in [20, Proposition 5.43]. We have that if $V,W$ are vector groups, and $\varphi:V\rightarrow W$ is a continuous homomorphism, then it is $\mathbb{R}$-linear and its image is a closed subspace of $W$ which is a direct summand [20, Theorem A2.12].